OpenStudy (anonymous):

Solve by completing the square. m^2-49/2m=1/2

OpenStudy (saifoo.khan):

\[\frac{2}{-49 - \sqrt{2409}} , \frac 14 (49+ \sqrt{2409})\]

OpenStudy (anonymous):

Can you show me the steps so I can compare them with mine?

OpenStudy (saifoo.khan):

Do u have to show them the steps as well? coz we can do other way as well.

OpenStudy (anonymous):

I just wanna see where I went wrong with it.

OpenStudy (saifoo.khan):

Ok, i solved using Quadratic formula.

OpenStudy (anonymous):

m^2-49/2m-1/2=0, so a=1 b=49/2 c=-1/2, to be used in the quadratic formula of -b+- square root of b^2-4ac/2a.

OpenStudy (saifoo.khan):

b = -49/2 And then \[\frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

Yes, right. But the directions tells me to complete the square.

OpenStudy (saifoo.khan):

that dosnt matter. u will get the same answer.

OpenStudy (anonymous):

Ok, I'm gonna try to work it out and see if I get what you gave me.

OpenStudy (anonymous):

m^2-49/2m=1/2 (m-49/4)^2-(49/4)^2=1/2 This should be the first step

OpenStudy (anonymous):

@andras, I get m^2-49/2m+2401/16, the 2401/16 in the rules for completing a square is take the coefficient of the second term divide by 2, then square it and add it to the end as the third term and also add it to the other side. So, 1/2 x 49/2= 49/4 ^2=2401/16. This is what I came up with after doing that (m^2-49/4)^2=2409/16?

OpenStudy (anonymous):

@saifoo.khan where are you?